OpenStudy (anonymous):

4.Form a quadratic equation whose roots are 2 times the roots of the quadratic equation 5x=2x^2-7 @ganeshie8 @Callisto @Jhannybean @zepdrix

OpenStudy (anonymous):

\[2x^2-5x-7=0\]

ganeshie8 (ganeshie8):

If a number "m" is a root of f(x) then we have f(m) = 0 Notice that the number "2m" is a root of f(x/2) because plugging in x = 2m gives you f(2m/2) = f(m) = 0

ganeshie8 (ganeshie8):

so simply replace "x" by "x/2" in the equation to get the equation whose roots are 2 times

OpenStudy (anonymous):

how u get \[f( \frac{ x }{ 2 })\] ?

ganeshie8 (ganeshie8):

replace every "x" by "x/2"

OpenStudy (anonymous):

okay, i undestand

ganeshie8 (ganeshie8):

\[f(x) = 2x^2-5x-7=0\]

ganeshie8 (ganeshie8):

the equation whose roots are 2 times the toots of f(x) would be : \[f(\frac{x}{2})~~ =~~~ 2(\frac{x}{2})^2-5(\frac{x}{2})-7=0\] simplify

ganeshie8 (ganeshie8):

2 needs to stay outside the square

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

sorry, please, if I call with x_1 and x_2 the roots of your equation, then I can write this: \[x _{1}+x _{2}=\frac{ 5 }{ 2 }\] and \[x _{1}*x _{2}=-\frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

\[2(\frac{ x^2 }{ 4 })-\frac{ 5x }{ 2 }-7=0\] like this? @ganeshie8

OpenStudy (michele_laino):

now in serching for an equation whose roots are 2 times x_1 and x_2, I write my new roots as below: \[z _{1}=2 x _{1},z _{2}=2*x _{2}\]

ganeshie8 (ganeshie8):

yes looks good, simplify

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ x^2 }{ 2 }-\frac{ 5x }{ 2 }-7=0\]

OpenStudy (anonymous):

\[x^2-5x-14=0\]

ganeshie8 (ganeshie8):

Perfect !

OpenStudy (anonymous):

thnx a lot @ganeshie8 and @Michele_Laino

OpenStudy (michele_laino):

Thank you! @MARC_

ganeshie8 (ganeshie8):

In general : the equation whose roots are "k" times the roots of f(x)=0 is given by f(x/k)=0

OpenStudy (anonymous):

okay,thnx again @ganeshie8